In 2 years I will be 5 times as old as my son and half as old as my father. My father just turned 78 . How old is my son?
(A) 5 (B) 6 (C) 7 (D) 8 (E) 9

Answer by richwmiller(17219)

x+2=5*(s+2)
x+2=(f+2)/2
f=78
s=6 x=38